AS13030

Status Info

Invalid ticket.

Current Outages

Duration:15.08.16 12:00 -
Severity (1-5): 1
Status:In Progress
Description:

We are currently experience reccuring IPv6 routing issues in our network. Some customers can't reach certain IP addresses or the internet at whole.

Investigation is in progress.

Update 18.08.2016 16:15 CEST

We are tracking this issue in #10250.

Update 22.08.2016 16:15 CEST

The Maintenance in #10250 didn't resolve the Routing Issue.

Investigation is still in progress.There is no ETR yet.

Update 25.08.2016 11:00 CEST

We are in contact with our vendor, unfortunately still no ETR.

Update 31.08.2016 17:30 CEST

We have a workaround in place and hope this stabilizes the situation for IPv6 connectivity. Debugging with our vendor is still ongoing.

Update 25.10.2016 15:20 CEST

The case has been escalated at our supplier, both administratively and on a technical level. It is beeing processed by their 3rd level protocol team.

Update 16.12.2016

The router replace according to ticket #10611 did not solve the problem. Further technical support from our supplier is needed.

Duration:14.02.17 16:07 -
Severity (1-5): 1
Status:In Progress
Description:

We are experiencing an outage of our links between Munich, SpaceNet and Amsterdam, Nikhef as well as Munich, SpaceNet and London, Telehouse East as well as Zurich, Equinix and Frankfurt, Interxion. Because of our redundant network design, we do not expect any loss of service. Traffic is routed via alternate paths.

Update 16:52 CET

Supplier has located the problem in in Frankfurt, Germany. Technicians are on the way.

Update 19:28 CET

The fault location has been identified at a construction site in Frankfurt. It is in the area of Europa-Allee and Güterplatz. A cable with 864 fibres has been damaged due to drilling. Fibre crews are currently attempting to pull the cable back to the next available chamber.  At the moment it is also planned to bring a replacement cable on site in case a temporary solution might become necessary. No ETR is available at present as damaged assessment is still required.

Update 20:53 CET

All available fibre engineers of our supplier have been activated and are on site. The primary action plan is still to attempt to pull the cable back to the next available chamber that has not been damaged. If that fails, the secondary action plan will be that the local fire department builds a cable ladder over the street to install all cables temporary. At the very moment we are following the primary action plan and preparation works are ongoing. As stated, if this fails we run the cables over the new cable ladder. We have been provided with an current ETR of 16 hours.

Update 23:05 CET

The fibre engineers have found an undamaged duct. The cable have been cut already. 
Pulling the cable into the undamaged duct will start shortly. Works are ongoing. Current ETR is 14hours.

Update 2017-02-15 02:39 CET

Repair works are currently ongoing. We have received an update that, due to the coordination with other companies on site, they've some difficulties to pull the cables. this will most likely take an additional 3-4 hrs from now. An updated ETR will be sent once splicing begins.

Update 2017-02-15 06:23 CET

The replacement of the broken fibre has been completed. Preparations are being made to begin fibre splicing

Update 2017-02-15 07:46 CET

We have been advised by the supplier that all fibre technicians are onsite.
Following actions have been taken so far:
Duct system has been checked and one single tube was found to be not damaged
500 meters of a new 864 fibre cable was installed
600 meters of a new 192 fibre cable was installed
The fibre crew is currently installing the cable now into the new joints.
Splicing will commence at approx 11 am, services will be restored as works progress but due to the size of the cable the completion of the works is yet to be determined
Update 08:10 CET

The link between Frankfurt, Interxion and Zurich, Equinix is back up since 07:53

Update 15:05 CET
We have been advised that the splicing is progressing but not finished. However, we see all our links back up again.
Duration:04.02.17 19:35 -
Severity (1-5): 2
Status:In Progress
Description:

We have been experiencing a DDOS attack against an Init7 customer on 4.2.2017 19:35 CET. After having taken countermeasures we see the situation stable.

For further information please take a look at:

https://drive.google.com/file/d/0B1dniAyu5OGWVzAxa24yZnJrdHM/view

Duration:01.02.17 10:40 -
Severity (1-5): 2
Status:In Progress
Description:

We have experienced an unexpected supvisor reboot on our Fiber7 POP Switch 790ENG.

The module is up again and we see all services back online.

Investigation for the cause is in progress.

Upcoming Maintenance

Duration:16.03.17 01:00 - 16.03.17 09:00
Severity (1-5): 1
Status:Planned
Description:

Our supplier is conducting maintenance work on our link between Vienna, Interxion and London, Telehouse. Because of our redundant network design, we do not expect any loss of service. Traffic is routed via alternate paths.

Duration:09.03.17 04:00 - 09.03.17 07:00
Severity (1-5): 1
Status:Planned
Description:

Our supplier is conducting maintenance work on our link between London, Telehouse North (LON-2) and Paris, Telehouse Voltaire (PAR-1).

Because of our redundant network design, we do not expect any loss of service. Traffic is routed via alternate paths.